OpenStudy (gauge):

Find the surface area of the rectangular prism with l = 15.5 m, w = 16.5 m, and h = 4.5 m. (1 point)

OpenStudy (anonymous):

Surface Area (S.A.) \[S.A.= 2( wh + lw + lh) \] \[S.A. = [(16.5\times4.5)+(15.5\times16.5)+(15.5\times4.5)] \] \[S.A. = (74.25)+(255.75)+(69.75)\] \[S.A. = 399.75 \]